# DialectBench: A NLP Benchmark for Dialects, Varieties, and Closely-Related Languages

### v1

@@ C.1 DialectBench Variety list
Glottocode mBERT seen UDP POS NER SDQA RCMC NLI TC SC DI MT
albanian
albanian
alba1267 - - - - - - - -
gheg albanian
gheg1238 - - - - - - - - -
anglic
african american vernacular english
afri1276 - - - - - - - - - -
australian english
aust1314 - - - - - - - - - -
english
stan1293 - - -
english (a:scotland)
stan1293 - - - - - - - - - -
english (a:uk)
stan1293 - - - - - - - - - -
english (o:controlled)
stan1293 - - - - - - - - - -
indian english (a:north)
indi1255 - - - - - - - - - -
indian english (a:south)
indi1255 - - - - - - - - - -
irish english
iris1255 - - - - - - - - - -
jamaican creole english
jama1262 - - - - - - - - - -
kenyan english
keny1281 - - - - - - - - - -
new zealand english
newz1240 - - - - - - - - - -
nigerian english
nige1260 - - - - - - - - - -
north american english
nort3314 - - - - - - - - - -
old english (ca. 450-1100)
olde1238 - - - - - - - - - -
philippine english
phil1246 - - - - - - - - - -
singlish
sing1272 - - - - - - - - -
southeast american english
sout3300 - - - - - - - - - -
southern african english
sout3331 - - - - - - - - - -
arabic
aleppo
alep1241 - - - - - - - - -
algerian arabic
alge1239 - - - - - - -
arabian peninsula arabic (a:yemen)
arab1393 - - - - - - - - -
arabic (a:bahrain)
arabic (a:jordan)
arabic (a:saudi-arabia)
egyptian arabic
egyp1253 - - - -
egyptian arabic (a:alx)
egyp1253 - - - - - - - - -
egyptian arabic (a:asw)
egyp1253 - - - - - - - - -
egyptian arabic (a:cai)
egyp1253 - - - - - - - - -
egyptian arabic (a:kha)
egyp1253 - - - - - - - - -
fez. meknes
fezm1238 - - - - - - - - -
gilit mesopotamian arabic
meso1252 - - - - - - - - -
gulf arabic
gulf1241 - - - - - - - - - -
gulf arabic (a:doh)
gulf1241 - - - - - - - - -
gulf arabic (a:jed)
gulf1241 - - - - - - - - -
gulf arabic (a:mus)
gulf1241 - - - - - - - - -
gulf arabic (a:riy)
gulf1241 - - - - - - - - -
levantine arabic
nort3139 - - - - - - - - - -
levantine arabic (a:north)
nort3139 - - - - - - - -
levantine arabic (a:north-dam)
nort3139 - - - - - - - - -
libyan arabic
liby1240 - - - - - - - - - - -
libyan arabic (a:ben)
liby1240 - - - - - - - - -
moroccan arabic
moro1292 - - - - - -
najdi arabic
najd1235 - - - - - - - -
north african arabic
nort3191 - - - - - - - - -
north mesopotamian arabic
nort3142 - - - - - - - -
north mesopotamian arabic (a:bas)
nort3142 - - - - - - - - -
north mesopotamian arabic (a:mos)
nort3142 - - - - - - - - -
rabat-casablanca arabic
raba1252 - - - - - - - - -
sfax
sfax1238 - - - - - - - - -
south levantine arabic
sout3123 - - - - - -
south levantine arabic (a:south-amm)
sout3123 - - - - - - - - -
south levantine arabic (a:south-jer)
sout3123 - - - - - - - - -
south levantine arabic (a:south-sal)
sout3123 - - - - - - - - -
standard arabic
stan1318 - -
sunni beiruti arabic
sunn1238 - - - - - - - - -
ta’izzi-adeni arabic
taiz1242 - - - - - - - - -
tripolitanian arabic
trip1239 - - - - - - - - -
tunisian arabic
tuni1259 - - - - - - -
tunisian arabic (a:tun)
tuni1259 - - - - - - - - -
tunisian arabic (r:casual)
tuni1259 - - - - - - - - - -
basque
basque (a:amenduze)
basq1248 - - - - - - - - -
basque (a:azkaine)
basq1248 - - - - - - - - - -
basque (a:baigorri)
basq1248 - - - - - - - - - -
basque (a:barkoxe)
basq1248 - - - - - - - - - -
basque (a:donibane)
basq1248 - - - - - - - - - -
basque (a:garruze)
basq1248 - - - - - - - - - -
basque (a:iholdi)
basq1248 - - - - - - - - - -
basque (a:jatsu)
basq1248 - - - - - - - - - -
basque (a:jutsi)
basq1248 - - - - - - - - - -
basque (a:larzabale)
basq1248 - - - - - - - - - -
basque (a:luhuso)
basq1248 - - - - - - - - - -
basque (a:sara)
basq1248 - - - - - - - - - -
basque (a:senpere)
basq1248 - - - - - - - - - -
basque (a:suhuskune)
basq1248 - - - - - - - - - -
basque (a:uharte)
basq1248 - - - - - - - - - -
navarro-labourdin basque (a:behorlegi)
basq1249 - - - - - - - - - -
navarro-labourdin basque (a:bidarrai)
basq1249 - - - - - - - - - -
navarro-labourdin basque (a:helete)
basq1249 - - - - - - - - - -
navarro-labourdin basque (a:mugerre)
basq1249 - - - - - - - - - -
navarro-labourdin basque (a:urruna)
basq1249 - - - - - - - - - -
souletin (a:maule)
basq1250 - - - - - - - - - -
bengali
vanga (a:barisal)
vang1242 - - - - - - - - - -
vanga (a:dhaka)
vang1242 - - - - - - - -
vanga (a:jessore)
vang1242 - - - - - - - - - -
vanga (a:khulna)
vang1242 - - - - - - - - - -
vanga (a:kushtia)
vang1242 - - - - - - - - - -
vanga (a:west bengal)
vang1242 - - - - - - - - -
circassian
adyghe
adyg1241 - - - - - - - - - -
kabardian
kaba1278 - - - - - - - - - -
common turkic
bashkir
bash1264 - - - - - - - - -
central oghuz
azer1255 - - - - - - - - - -
central oghuz (m:spoken)
azer1255 - - - - - - - - -
crimean tatar
crim1257 - - - - - - - - - -
kara-kalpak
kara1467 - - - - - - - - - -
kazakh
kaza1248 - - - - - - - - -
kirghiz
kirg1245 - - - - - - - - -
north azerbaijani
nort2697 - - - - - - -
sakha
yaku1245 - - - - - - - - - -
south azerbaijani
sout2697 - - - - - - -
turkish
nucl1301 - - - - - - - -
uzbek
uzbe1247 - - - - - - - - -
eastern romance
aromanian
arom1237 - - - - - - - - - -
moldavian
mold1248 - - - - - - - - - -
romanian
roma1327 - - - - - - - - -
eastern-western armenian
eastern armenian
nucl1235 - - - - - - - -
western armenian
homs1234 - - - - - - - - -
farsic
dari
dari1249 - - - - - - - - - -
frisian
ems-weser frisian
sate1242 - - - - - - - - - -
northern frisian
nort2626 - - - - - - - - - -
western frisian
west2354 - - - - - - - - -
gallo-italian
emiliano-romagnolo
emil1243 - - - - - - - - - -
ligurian
ligu1248 - - - - - -
lombard
lomb1257 - - - - - - -
piemontese
piem1238 - - - - - - - - -
venetian
vene1258 - - - - - - - -
gallo-rhaetian
anglo-norman
angl1258 - - - - - - - - - -
arpitan
fran1260 - - - - - - - - - -
french
stan1290 - - - - - - -
french (a:paris)
stan1290 - - - - - - - - -
friulian
friu1240 - - - - - - - -
old french (842-ca. 1400)
oldf1239 - - - - - - - - -
romansh
roma1326 - - - - - - - - - -
walloon
wall1255 - - - - - - - - - -
greater panjabic
eastern panjabi
panj1256 - - - - - - - - -
western panjabi
west2386 - - - - - - - - -
greek
apulian greek
apul1237 - - - - - - - - - -
cretan
cret1244 - - - - - - - - - -
cypriot greek
cypr1249 - - - - - - - - - - -
cypriot greek (r:casual, m:written, i:fb)
cypr1249 - - - - - - - - - -
cypriot greek (r:casual, m:written, i:other)
cypr1249 - - - - - - - - - -
cypriot greek (r:casual, m:written, i:twitter)
cypr1249 - - - - - - - - - -
modern greek
mode1248 - - - - - - - - -
modern greek (r:casual, m:written, i:fb)
mode1248 - - - - - - - - - -
modern greek (r:casual, m:written, i:other)
mode1248 - - - - - - - - - -
modern greek (r:casual, m:written, i:twitter)
mode1248 - - - - - - - - - -
pontic
pont1253 - - - - - - - - - -
high german
bavarian
bava1246 - - - - - - - - -
central alemannic
swis1247 - - - - - - - - - -
central alemannic (a:ag)
swis1247 - - - - - - - - - -
central alemannic (a:ai)
swis1247 - - - - - - - - - -
central alemannic (a:ar)
swis1247 - - - - - - - - - -
central alemannic (a:be)
swis1247 - - - - - - - - -
central alemannic (a:bl)
swis1247 - - - - - - - - - -
central alemannic (a:bs)
swis1247 - - - - - - - - -
central alemannic (a:fr)
swis1247 - - - - - - - - - -
central alemannic (a:gl)
swis1247 - - - - - - - - - -
central alemannic (a:gr)
swis1247 - - - - - - - - - -
central alemannic (a:lu)
swis1247 - - - - - - - - -
central alemannic (a:nw)
swis1247 - - - - - - - - - -
central alemannic (a:ow)
swis1247 - - - - - - - - - -
central alemannic (a:sg)
swis1247 - - - - - - - - - -
central alemannic (a:sh)
swis1247 - - - - - - - - - -
central alemannic (a:so)
swis1247 - - - - - - - - - -
central alemannic (a:sz)
swis1247 - - - - - - - - - -
central alemannic (a:tg)
swis1247 - - - - - - - - - -
central alemannic (a:ur)
swis1247 - - - - - - - - - -
central alemannic (a:vs)
swis1247 - - - - - - - - - -
central alemannic (a:zg)
swis1247 - - - - - - - - - -
central alemannic (a:zh)
swis1247 - - - - - - -
central bavarian
cent1967 - - - - - - - - - -
german
stan1295 - - - - - - -
kölsch
kols1241 - - - - - - - - - -
limburgan
limb1263 - - - - - - - -
luxemburgish
luxe1243 - - - - - - -
pennsylvania german
penn1240 - - - - - - - - - -
pfaelzisch-lothringisch
pala1330 - - - - - - - - - -
upper saxon
uppe1465 - - - - - - - - - -
hindustani
fiji hindi
fiji1242 - - - - - - - - - -
hindi
hind1269 - - - - - - - - -
inuit
alaskan inupiaq
inup1234 - - - - - - - - - -
kalaallisut
kala1399 - - - - - - - - - -
italian romance
continental southern italian
neap1235 - - - - - - - -
italian
ital1282 - - - - -
italian (a:abruzzo)
ital1282 - - - - - - - - - -
italian (a:basilicata)
ital1282 - - - - - - - - - -
italian (a:calabria)
ital1282 - - - - - - - - - -
italian (a:campania)
ital1282 - - - - - - - - - -
italian (a:emilia-romagna)
ital1282 - - - - - - - - - -
italian (a:friuli-venezia giulia)
ital1282 - - - - - - - - - -
italian (a:lazio)
ital1282 - - - - - - - - - -
italian (a:liguria)
ital1282 - - - - - - - - - -
italian (a:lombardia)
ital1282 - - - - - - - - - -
italian (a:marche)
ital1282 - - - - - - - - - -
italian (a:molise)
ital1282 - - - - - - - - - -
italian (a:piemonte)
ital1282 - - - - - - - - - -
italian (a:puglia)
ital1282 - - - - - - - - - -
italian (a:sardegna)
ital1282 - - - - - - - - - -
italian (a:sicilia)
ital1282 - - - - - - - - - -
italian (a:toscana)
ital1282 - - - - - - - - - -
italian (a:trentino-alto adige/südtirol)
ital1282 - - - - - - - - - -
italian (a:umbria)
ital1282 - - - - - - - - - -
italian (a:unknown)
ital1282 - - - - - - - - - -
italian (a:veneto)
ital1282 - - - - - - - - - -
italian (r:casual, m:written, i:tweet)
ital1282 - - - - - - - -
italian (r:formal, m:written, i:essay)
ital1282 - - - - - - - -
italian romance (a:apulia, m:spoken, i:tarantino)
sicilian
sici1248 - - - - - - -
komi
komi
komi1267 - - - - - - - - - -
komi-permyak
komi1269 - - - - - - - -
komi-zyrian (m:spoken)
komi1268 - - - - - - - - -
komi-zyrian (m:written)
komi1268 - - - - - - - - -
korean
korean (a:south-eastern, m:spoken)
kore1280 - - - - - - - - - -
seoul (m:spoken)
seou1239 - - - - - - - - -
kurdish
central kurdish
cent1972 - - - - - - - -
kurdish
kurd1259 - - - - - - - - - -
northern kurdish
nort2641 - - - - - - - - -
sine’i
sine1239 - - - - - - - - - -
sorani
sora1257 - - - - - - - - - -
latvian
east latvian
east2282 - - - - - - - -
latvian
latv1249 - - - - - - -
mari
eastern mari
east2328 - - - - - - - - - -
western mari
west2392 - - - - - - - - - -
modern dutch
dutch
dutc1256 - - - - - - -
western flemish
vlaa1240 - - - - - - - - - -
zeeuws
zeeu1238 - - - - - - - - - -
neva
central and north pohjanmaa (a:ostrobothnian)
cent1985 - - - - - - - - - -
estonian
esto1258 - - - - - - - - -
finnish
finn1318 - - - - - - - - -
häme (a:tavastian)
hame1240 - - - - - - - - - -
neva (a:south-west trans)
savo (a:savonian)
savo1254 - - - - - - - - - -
southeastern finnish (a:south-east)
sout1743 - - - - - - - - - -
southwestern finnish (a:south-west)
sout2677 - - - - - - - - - -
norwegian
norwegian
norw1258 - - - - - - - - - -
norwegian (a:eastern)
norw1258 - - - - - - - - - -
norwegian (a:setesdal)
norw1258 - - - - - - - - - -
norwegian (a:southwestern)
norw1258 - - - - - - - - - -
norwegian (m:written, i:samnorsk)
norwegian bokmål (m:written)
norw1259 - - - - -
norwegian nynorsk (m:written)
norw1262 - - - - -
norwegian nynorsk (m:written, i:old)
norw1262 - - - - - - - - -
saami
north saami
nort2671 - - - - - - - - -
skolt saami
skol1241 - - - - - - - - -
sabellic
umbrian
umbr1253 - - - - - - - - -
sardo-corsican
corsican
cors1241 - - - - - - - - - -
sardinian
sard1257 - - - - - - - -
serbian-croatian-bosnian
bosnian standard
bosn1245 - - - - - - - - -
croatian standard
croa1245 - - - - - - - - -
serbian standard
serb1264 - - - - - - - - -
serbian-croatian-bosnian
sout1528 - - - - - - - - -
sinitic
cantonese
cant1236 - - - - - - - -
classical chinese
lite1248 - - - - - - - -
classical-middle-modern sinitic (a:hongkong, o:traditional)
clas1255 - - - - - - - - -
classical-middle-modern sinitic (o:simplified)
clas1255 - - - - -
classical-middle-modern sinitic (o:traditional)
clas1255 - - - - - - -
hakka chinese
hakk1236 - - - - - - - - - -
mandarin chinese
mand1415 - - - - - - - - - -
mandarin chinese (a:mainland, o:simplified)
mand1415 - - - - - - - - - -
mandarin chinese (a:mainland, o:traditional, i:synthetic)
mand1415 - - - - - - - - - -
mandarin chinese (a:taiwan, o:simplified)
mand1415 - - - - - - - - - -
mandarin chinese (a:taiwan, o:traditional, i:synthetic)
mand1415 - - - - - - - - - -
min nan chinese
minn1241 - - - - - - - - - -
wu chinese
wuch1236 - - - - - - - - - -
sorbian
lower sorbian
lowe1385 - - - - - - - - - -
upper sorbian
uppe1395 - - - - - - - - - -
sotho-tswana (s.30)
northern sotho
nort3233 - - - - - - -
southern sotho
sout2807 - - - - - - -
southwestern shifted romance
brazilian portuguese
braz1246 - - - - - - - -
extremaduran
extr1243 - - - - - - - - - -
galician
gali1258 - - - - - - -
latin american spanish
amer1254 - - - - - - - - - -
mirandese
mira1251 - - - - - - - - - -
occitan
occi1239 - - - - -
portuguese (a:european)
port1283 - - - - - -
portuguese (i:mix)
port1283 - - - - - - - - -
portuguese (m:written)
port1283 - - - - - - -
spanish
stan1288 - - - - - -
spanish (a:europe)
stan1288 - - - - - - - - - -
swahili
swahili
swah1253 - - - - - - - - - -
swahili (a:kenya)
swah1253 - - - - - - - - - -
swahili (a:tanzania)
swah1253 - - - - - - - - - -
tigrinya
tigrinya (a:eritrea)
tigr1271 - - - - - - - - - -
tigrinya (a:ethiopia)
tigr1271 - - - - - - - - - -
tupi-guarani subgroup i.a
mbyá guaraní (a:brazil)
mbya1239 - - - - - - - - -
mbyá guaraní (a:paraguay)
mbya1239 - - - - - - - - -
old guarani
oldp1258 - - - - - - - - -
west low german
west low german
west2357 - - - - - - -
yoruba
yoruba (a:central nigeria)
yoru1245 - - - - - - - - -

### v2

@@ C.1 DialectBench Variety list
Language cluster Variety name Glottocode mBERT seen UDP POS NER SDQA RCMC NLI TC SC DI MT
albanian albanian alba1267 - - - - - - - -
gheg albanian gheg1238 - - - - - - - - -
anglic african american vernacular english afri1276 - - - - - - - - - -
australian english aust1314 - - - - - - - - - -
english stan1293 - - -
english (a:scotland) stan1293 - - - - - - - - - -
english (a:uk) stan1293 - - - - - - - - - -
english (o:controlled) stan1293 - - - - - - - - - -
indian english (a:north) indi1255 - - - - - - - - - -
indian english (a:south) indi1255 - - - - - - - - - -
irish english iris1255 - - - - - - - - - -
jamaican creole english jama1262 - - - - - - - - - -
kenyan english keny1281 - - - - - - - - - -
new zealand english newz1240 - - - - - - - - - -
nigerian english nige1260 - - - - - - - - - -
north american english nort3314 - - - - - - - - - -
old english (ca. 450-1100) olde1238 - - - - - - - - - -
philippine english phil1246 - - - - - - - - - -
singlish sing1272 - - - - - - - - -
southeast american english sout3300 - - - - - - - - - -
southern african english sout3331 - - - - - - - - - -
arabic aleppo alep1241 - - - - - - - - -
algerian arabic alge1239 - - - - - - -
arabian peninsula arabic (a:yemen) arab1393 - - - - - - - - -
arabic (a:bahrain) - - - - - - - - - - -
arabic (a:jordan) - - - - - - - - - -
arabic (a:saudi-arabia) - - - - - - - - - -
egyptian arabic egyp1253 - - - -
egyptian arabic (a:alx) egyp1253 - - - - - - - - -
egyptian arabic (a:asw) egyp1253 - - - - - - - - -
egyptian arabic (a:cai) egyp1253 - - - - - - - - -
egyptian arabic (a:kha) egyp1253 - - - - - - - - -
fez. meknes fezm1238 - - - - - - - - -
gilit mesopotamian arabic meso1252 - - - - - - - - -
gulf arabic gulf1241 - - - - - - - - - -
gulf arabic (a:doh) gulf1241 - - - - - - - - -
gulf arabic (a:jed) gulf1241 - - - - - - - - -
gulf arabic (a:mus) gulf1241 - - - - - - - - -
gulf arabic (a:riy) gulf1241 - - - - - - - - -
levantine arabic nort3139 - - - - - - - - - -
levantine arabic (a:north) nort3139 - - - - - - - -
levantine arabic (a:north-dam) nort3139 - - - - - - - - -
libyan arabic liby1240 - - - - - - - - - - -
libyan arabic (a:ben) liby1240 - - - - - - - - -
moroccan arabic moro1292 - - - - - -
najdi arabic najd1235 - - - - - - - -
north african arabic nort3191 - - - - - - - - -
north mesopotamian arabic nort3142 - - - - - - - -
north mesopotamian arabic (a:bas) nort3142 - - - - - - - - -
north mesopotamian arabic (a:mos) nort3142 - - - - - - - - -
rabat-casablanca arabic raba1252 - - - - - - - - -
sfax sfax1238 - - - - - - - - -
south levantine arabic sout3123 - - - - - -
south levantine arabic (a:south-amm) sout3123 - - - - - - - - -
south levantine arabic (a:south-jer) sout3123 - - - - - - - - -
south levantine arabic (a:south-sal) sout3123 - - - - - - - - -
standard arabic stan1318 - -
sunni beiruti arabic sunn1238 - - - - - - - - -
ta’izzi-adeni arabic taiz1242 - - - - - - - - -
tripolitanian arabic trip1239 - - - - - - - - -
tunisian arabic tuni1259 - - - - - - -
tunisian arabic (a:tun) tuni1259 - - - - - - - - -
tunisian arabic (r:casual) tuni1259 - - - - - - - - - -
basque basque (a:amenduze) basq1248 - - - - - - - - -
basque (a:azkaine) basq1248 - - - - - - - - - -
basque (a:baigorri) basq1248 - - - - - - - - - -
basque (a:barkoxe) basq1248 - - - - - - - - - -
basque (a:donibane) basq1248 - - - - - - - - - -
basque (a:garruze) basq1248 - - - - - - - - - -
basque (a:iholdi) basq1248 - - - - - - - - - -
basque (a:jatsu) basq1248 - - - - - - - - - -
basque (a:jutsi) basq1248 - - - - - - - - - -
basque (a:larzabale) basq1248 - - - - - - - - - -
basque (a:luhuso) basq1248 - - - - - - - - - -
basque (a:sara) basq1248 - - - - - - - - - -
basque (a:senpere) basq1248 - - - - - - - - - -
basque (a:suhuskune) basq1248 - - - - - - - - - -
basque (a:uharte) basq1248 - - - - - - - - - -
navarro-labourdin basque (a:behorlegi) basq1249 - - - - - - - - - -
navarro-labourdin basque (a:bidarrai) basq1249 - - - - - - - - - -
navarro-labourdin basque (a:helete) basq1249 - - - - - - - - - -
navarro-labourdin basque (a:mugerre) basq1249 - - - - - - - - - -
navarro-labourdin basque (a:urruna) basq1249 - - - - - - - - - -
souletin (a:maule) basq1250 - - - - - - - - - -
bengali vanga (a:barisal) vang1242 - - - - - - - - - -
vanga (a:dhaka) vang1242 - - - - - - - -
vanga (a:jessore) vang1242 - - - - - - - - - -
vanga (a:khulna) vang1242 - - - - - - - - - -
vanga (a:kushtia) vang1242 - - - - - - - - - -
vanga (a:west bengal) vang1242 - - - - - - - - -
circassian adyghe adyg1241 - - - - - - - - - -
kabardian kaba1278 - - - - - - - - - -
common turkic bashkir bash1264 - - - - - - - - -
central oghuz azer1255 - - - - - - - - - -
central oghuz (m:spoken) azer1255 - - - - - - - - -
crimean tatar crim1257 - - - - - - - - - -
kara-kalpak kara1467 - - - - - - - - - -
kazakh kaza1248 - - - - - - - - -
kirghiz kirg1245 - - - - - - - - -
north azerbaijani nort2697 - - - - - - -
sakha yaku1245 - - - - - - - - - -
south azerbaijani sout2697 - - - - - - -
turkish nucl1301 - - - - - - - -
uzbek uzbe1247 - - - - - - - - -
eastern romance aromanian arom1237 - - - - - - - - - -
moldavian mold1248 - - - - - - - - - -
romanian roma1327 - - - - - - - - -
eastern-western armenian eastern armenian nucl1235 - - - - - - - -
western armenian homs1234 - - - - - - - - -
farsic dari dari1249 - - - - - - - - - -
frisian ems-weser frisian sate1242 - - - - - - - - - -
northern frisian nort2626 - - - - - - - - - -
western frisian west2354 - - - - - - - - -
gallo-italian emiliano-romagnolo emil1243 - - - - - - - - - -
ligurian ligu1248 - - - - - -
lombard lomb1257 - - - - - - -
piemontese piem1238 - - - - - - - - -
venetian vene1258 - - - - - - - -
gallo-rhaetian anglo-norman angl1258 - - - - - - - - - -
arpitan fran1260 - - - - - - - - - -
french stan1290 - - - - - - -
french (a:paris) stan1290 - - - - - - - - -
friulian friu1240 - - - - - - - -
old french (842-ca. 1400) oldf1239 - - - - - - - - -
romansh roma1326 - - - - - - - - - -
walloon wall1255 - - - - - - - - - -
greater panjabic eastern panjabi panj1256 - - - - - - - - -
western panjabi west2386 - - - - - - - - -
greek apulian greek apul1237 - - - - - - - - - -
cretan cret1244 - - - - - - - - - -
cypriot greek cypr1249 - - - - - - - - - - -
cypriot greek (r:casual, m:written, i:fb) cypr1249 - - - - - - - - - -
cypriot greek (r:casual, m:written, i:other) cypr1249 - - - - - - - - - -
cypriot greek (r:casual, m:written, i:twitter) cypr1249 - - - - - - - - - -
modern greek mode1248 - - - - - - - - -
modern greek (r:casual, m:written, i:fb) mode1248 - - - - - - - - - -
modern greek (r:casual, m:written, i:other) mode1248 - - - - - - - - - -
modern greek (r:casual, m:written, i:twitter) mode1248 - - - - - - - - - -
pontic pont1253 - - - - - - - - - -
high german bavarian bava1246 - - - - - - - - -
central alemannic swis1247 - - - - - - - - - -
central alemannic (a:ag) swis1247 - - - - - - - - - -
central alemannic (a:ai) swis1247 - - - - - - - - - -
central alemannic (a:ar) swis1247 - - - - - - - - - -
central alemannic (a:be) swis1247 - - - - - - - - -
central alemannic (a:bl) swis1247 - - - - - - - - - -
central alemannic (a:bs) swis1247 - - - - - - - - -
central alemannic (a:fr) swis1247 - - - - - - - - - -
central alemannic (a:gl) swis1247 - - - - - - - - - -
central alemannic (a:gr) swis1247 - - - - - - - - - -
central alemannic (a:lu) swis1247 - - - - - - - - -
central alemannic (a:nw) swis1247 - - - - - - - - - -
central alemannic (a:ow) swis1247 - - - - - - - - - -
central alemannic (a:sg) swis1247 - - - - - - - - - -
central alemannic (a:sh) swis1247 - - - - - - - - - -
central alemannic (a:so) swis1247 - - - - - - - - - -
central alemannic (a:sz) swis1247 - - - - - - - - - -
central alemannic (a:tg) swis1247 - - - - - - - - - -
central alemannic (a:ur) swis1247 - - - - - - - - - -
central alemannic (a:vs) swis1247 - - - - - - - - - -
central alemannic (a:zg) swis1247 - - - - - - - - - -
central alemannic (a:zh) swis1247 - - - - - - -
central bavarian cent1967 - - - - - - - - - -
german stan1295 - - - - - - -
kölsch kols1241 - - - - - - - - - -
limburgan limb1263 - - - - - - - -
luxemburgish luxe1243 - - - - - - -
pennsylvania german penn1240 - - - - - - - - - -
pfaelzisch-lothringisch pala1330 - - - - - - - - - -
upper saxon uppe1465 - - - - - - - - - -
hindustani fiji hindi fiji1242 - - - - - - - - - -
hindi hind1269 - - - - - - - - -
inuit alaskan inupiaq inup1234 - - - - - - - - - -
kalaallisut kala1399 - - - - - - - - - -
italian romance continental southern italian neap1235 - - - - - - - -
italian ital1282 - - - - -
italian (a:abruzzo) ital1282 - - - - - - - - - -
italian (a:basilicata) ital1282 - - - - - - - - - -
italian (a:calabria) ital1282 - - - - - - - - - -
italian (a:campania) ital1282 - - - - - - - - - -
italian (a:emilia-romagna) ital1282 - - - - - - - - - -
italian (a:friuli-venezia giulia) ital1282 - - - - - - - - - -
italian (a:lazio) ital1282 - - - - - - - - - -
italian (a:liguria) ital1282 - - - - - - - - - -
italian (a:lombardia) ital1282 - - - - - - - - - -
italian (a:marche) ital1282 - - - - - - - - - -
italian (a:molise) ital1282 - - - - - - - - - -
italian (a:piemonte) ital1282 - - - - - - - - - -
italian (a:puglia) ital1282 - - - - - - - - - -
italian (a:sardegna) ital1282 - - - - - - - - - -
italian (a:sicilia) ital1282 - - - - - - - - - -
italian (a:toscana) ital1282 - - - - - - - - - -
italian (a:trentino-alto adige/südtirol) ital1282 - - - - - - - - - -
italian (a:umbria) ital1282 - - - - - - - - - -
italian (a:unknown) ital1282 - - - - - - - - - -
italian (a:veneto) ital1282 - - - - - - - - - -
italian (r:casual, m:written, i:tweet) ital1282 - - - - - - - -
italian (r:formal, m:written, i:essay) ital1282 - - - - - - - -
italian romance (a:apulia, m:spoken, i:tarantino) - - - - - - - - - - -
sicilian sici1248 - - - - - - -
komi komi komi1267 - - - - - - - - - -
komi-permyak komi1269 - - - - - - - -
komi-zyrian (m:spoken) komi1268 - - - - - - - - -
komi-zyrian (m:written) komi1268 - - - - - - - - -
korean korean (a:south-eastern, m:spoken) kore1280 - - - - - - - - - -
seoul (m:spoken) seou1239 - - - - - - - - -
kurdish central kurdish cent1972 - - - - - - - -
kurdish kurd1259 - - - - - - - - - -
northern kurdish nort2641 - - - - - - - - -
sine’i sine1239 - - - - - - - - - -
sorani sora1257 - - - - - - - - - -
latvian east latvian east2282 - - - - - - - -
latvian latv1249 - - - - - - -
mari eastern mari east2328 - - - - - - - - - -
western mari west2392 - - - - - - - - - -
modern dutch dutch dutc1256 - - - - - - -
western flemish vlaa1240 - - - - - - - - - -
zeeuws zeeu1238 - - - - - - - - - -
neva central and north pohjanmaa (a:ostrobothnian) cent1985 - - - - - - - - - -
estonian esto1258 - - - - - - - - -
finnish finn1318 - - - - - - - - -
häme (a:tavastian) hame1240 - - - - - - - - - -
neva (a:south-west trans) - - - - - - - - - - -
savo (a:savonian) savo1254 - - - - - - - - - -
southeastern finnish (a:south-east) sout1743 - - - - - - - - - -
southwestern finnish (a:south-west) sout2677 - - - - - - - - - -
norwegian norwegian norw1258 - - - - - - - - - -
norwegian (a:eastern) norw1258 - - - - - - - - - -
norwegian (a:setesdal) norw1258 - - - - - - - - - -
norwegian (a:southwestern) norw1258 - - - - - - - - - -
norwegian (m:written, i:samnorsk) - - - - - - - - - - -
norwegian bokmål (m:written) norw1259 - - - - -
norwegian nynorsk (m:written) norw1262 - - - - -
norwegian nynorsk (m:written, i:old) norw1262 - - - - - - - - -
saami north saami nort2671 - - - - - - - - -
skolt saami skol1241 - - - - - - - - -
sabellic umbrian umbr1253 - - - - - - - - -
sardo-corsican corsican cors1241 - - - - - - - - - -
sardinian sard1257 - - - - - - - -
serbian-croatian-bosnian bosnian standard bosn1245 - - - - - - - - -
croatian standard croa1245 - - - - - - - - -
serbian standard serb1264 - - - - - - - - -
serbian-croatian-bosnian sout1528 - - - - - - - - -
sinitic cantonese cant1236 - - - - - - - -
classical chinese lite1248 - - - - - - - -
classical-middle-modern sinitic (a:hongkong, o:traditional) clas1255 - - - - - - - - -
classical-middle-modern sinitic (o:simplified) clas1255 - - - - -
classical-middle-modern sinitic (o:traditional) clas1255 - - - - - - -
hakka chinese hakk1236 - - - - - - - - - -
mandarin chinese mand1415 - - - - - - - - - -
mandarin chinese (a:mainland, o:simplified) mand1415 - - - - - - - - - -
mandarin chinese (a:mainland, o:traditional, i:synthetic) mand1415 - - - - - - - - - -
mandarin chinese (a:taiwan, o:simplified) mand1415 - - - - - - - - - -
mandarin chinese (a:taiwan, o:traditional, i:synthetic) mand1415 - - - - - - - - - -
min nan chinese minn1241 - - - - - - - - - -
wu chinese wuch1236 - - - - - - - - - -
sorbian lower sorbian lowe1385 - - - - - - - - - -
upper sorbian uppe1395 - - - - - - - - - -
sotho-tswana (s.30) northern sotho nort3233 - - - - - - -
southern sotho sout2807 - - - - - - -
southwestern shifted brazilian portuguese braz1246 - - - - - - - -
romance extremaduran extr1243 - - - - - - - - - -
galician gali1258 - - - - - - -
latin american spanish amer1254 - - - - - - - - - -
mirandese mira1251 - - - - - - - - - -
occitan occi1239 - - - - -
portuguese (a:european) port1283 - - - - - -
portuguese (i:mix) port1283 - - - - - - - - -
portuguese (m:written) port1283 - - - - - - -
spanish stan1288 - - - - - -
spanish (a:europe) stan1288 - - - - - - - - - -
swahili swahili swah1253 - - - - - - - - - -
swahili (a:kenya) swah1253 - - - - - - - - - -
swahili (a:tanzania) swah1253 - - - - - - - - - -
tigrinya tigrinya (a:eritrea) tigr1271 - - - - - - - - - -
tigrinya (a:ethiopia) tigr1271 - - - - - - - - - -
tupi-guarani subgroup i.a mbyá guaraní (a:brazil) mbya1239 - - - - - - - - -
mbyá guaraní (a:paraguay) mbya1239 - - - - - - - - -
old guarani oldp1258 - - - - - - - - -
west low german west low german west2357 - - - - - - -
yoruba yoruba (a:central nigeria) yoru1245 - - - - - - - - -
Table 8: Varieties represented in DialectBench.